OpenStudy (anonymous):

You invest $4,000.00 in a stock. The stock plan increases 8% the first year, decreases in value by 4% the second year, and increases by 6% the third year. I'm just not sure how to calculate the increases and decreases.

OpenStudy (anonymous):

Well first of all you start off with $4,000.00. Then it increases by 8% . So in order to calculate how much it has increased it would look like this. 4000 * 0.08 = ____ the amount it has increased.

OpenStudy (anonymous):

Then you add the amount it has increased to the 4000. Then follow the same process again with the second and third year.

OpenStudy (anonymous):

thank you! the increase then decrease threw me off for some reason i thought i had to do something different with the decrease of 4% . but thank you again!!

OpenStudy (anonymous):

Example: We start with 1000. So the first one increases in value 3% . So we have 1030 now(1000*0.03). The second decreases in value by 2% so 1030 * 0.02= 20.6. This means 1030 - 20.6 = 1009.4. Then on the third it increases by 2%. So this time 1009.4*0.02 = 20.188. So you end up with 1009.4 + 20.188 = $1029.588